OpenStudy (anonymous):

what is the least positive integer for which 12x is the cube?

Parth (parthkohli):

Hmm. Here's a list:\[1,8,27,64,125,216,343\cdots\]

Parth (parthkohli):

A list of cubes.

OpenStudy (anonymous):

12 = 3*4 = 3 * 2 * 2 So..To Make it Perfect Cube We have to Get Like this 2*2*2 * 3*3*3 So..as u See wat is Remaining 3 * 3 * 2 = 18

OpenStudy (anonymous):

This Is One of the Appraoch....I Suggest U To go By Parth 's Method..as It iS Easy

OpenStudy (anonymous):

12x = 216 x =18